5.0 kg of hydrogen is reacted with 31 kg of oxygen. The experiment yields 28 kg of water. Calculate the percent yield of the reaction


1
Expert's answer
2021-10-30T11:15:55-0400

2H2 + O2 → 2H2O

M(H2) = 2 g/mol

n(H2) "= \frac{5000}{2} = 2500 \;mol"

M(O2) = 32 g/mol

n(O2) "= \frac{31000}{32} = 968.75 \;mol"

According to the reaction for one mole of O2 we need two moles of H2. So, O2 is limiting reactant.

n(H2O) = 2n(O2) "= 2 \times 968.75 = 1937.5 \;mol"

M(H2O) = 18 g/mol

m(H2O)(theoretical) "= 1937.5 \times 18 = 34875 \;g = 34.875 \;kg"

Proportion:

37.875 kg – 100 %

28 kg – x

"x = \frac{28 \times 100}{37.875} = 73.927 \; \%"

Answer: 74 %